A manufacturing plant conducted a survey to determine it's employees reaction toward a proposed change in working hours.a breakdown of the response is as follows: production:17 agree 23 disagree. Office:8 agree 2 disagree. Suppose an employee is chosen at random with events being defined as A:the employee works in the production team. B:the employee agreed with the proposed change. C: the employee works in the office. D: the employee disagree with the proposed change. 4.2.1) P(A). 4.2.2) P(A or B). 4.2.3)P(C and D). 4.2.4)P(c/D)
1
Expert's answer
2020-04-16T19:05:16-0400

1) "P(A)=\frac{17+23}{17+23+8+2}=\frac{40}{50}=0.8."


2) "P(AorB)=\frac{17+23+8}{50}=0.96."


3) "P(CandD)=\frac{2}{50}=0.04."


4) "P(C|D)=\frac{2}{23+2}=0.08."